The expression 2x^3+x^2-7x-6 factorizes into a product of three factors.
Given that two of the factors are (x+1) and (x-2), write down its third factor.

Answer by lwsshak3(11628)

The expression 2x^3+x^2-7x-6 factorizes into a product of three factors.
Given that two of the factors are (x+1) and (x-2), write down its third factor.
**
use synthetic division:
.....2....|...2......1.......-7.....-6.......
.......................4........10.....6....
..............2.......5.........3......0...
(x-2)(2x^2+5x+3)=0
(x-2)(x+1)(2x+3)=0
third factor: (2x+3)